A Complete Bibliography of Publications in SIGCSE Bulletin: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

27 August 2024 Version 1.28

Title word cross-reference

12 [Wol11]. ’18 [PSTT17].

2012 [AS13b, Wal12a]. 2013 [Alv13, Bar13, Cam13, Car13, CCS13, Cut13, Edi14, Swe13]. 2014 [CDS13, DN14, ND14, Rod14, Tym14a]. 2015 [DeL14a, DeL14b, She15b]. 2016 [AT16a, CCV16, CCV17, MR15, SDCT16a, SDCT16c, She16a]. 2017 [DG17, EC17, Fit18b, MJ18, TCSM17, TCM^{+}17]. 2018 [BG18b, DD17, Dec19b, Gol18, JI18, MKMP18, MMK18, PQH17, She18]. 2019 [Cut19, FI19, HPQ19, MP19a, MPRM19, MP19b, MS19, Rod18, She19a]. 2020 [DH19].

42nd [C^{+}11].

50th [MS18, Set17a].

Academy [Had15]. access [Bar17a, Dec19a, LB18]. ACM [C^{+}11, Owe11b, ZCLR19, Fit15a, IS13, KD14, Owe12a, Owe12b, Set17a, Set17b, Set18a, Wal11a]. activities [Joy11]. add [Tow16]. Advancing [PT18]. against [Set18a]. ahead [CDS13]. Air [Had15]. Amy [Mil19]. anniversary [MS18]. Announcing [DWS19]. annual [McC12]. art [Wal10]. assessment [AO12]. Australasian [Sim13]. authorship [Set17b]. award [Alv13, DWS19, DW19, Dec19d, Doy16d]. awarded [Bal11, Bal12]. awardees [Set15]. awards [Bal13, Dec18b, Dec19b, Set14a, Set16a, She15a, She19a, Tym14c, Tym15, BF14, Tym14a].

Bard [Shu15]. based [AO12]. Berry [Ham15]. best [Alv13]. birthday [Ada11]. Board [Set16b, Dec19c, Rod19b, Owe11a, Rod19a]. BP [Bar15, KB15]. Broadening [BT15, Bar17a, Bar17b, PS16, RHM18, Tow16, BPG15, Bar15, BP16, PB16, PT18]. Bucks [Ano15a]. business [She14, She15b, She16b].

call [Cut19, FI19, MP19a, ZSH^{+}19]. Calling [FI19, JI18, MJ18, SM17]. Calvin [Ada11]. Camp [FP12]. Canterbury [Swe13]. categorization [Wal10]. CECC [Guz11]. celebrate [DEA18]. Celebrating [MS18]. Celebration [Owe12b, Set17a, Ada11]. Centenary [Owe12b]. chair [LGE12, McC12, Set17c, Set18b, Rod15b, Rod16, Set16b]. chairs [DeL14a, DeL14b, Mah18, Mah16]. change [HZPQH18a]. Changes [TY13a]. Chapman [Ano17b]. chapter [Rod17, Sim13]. Charlotte [Bar12]. Chengdu [Rod18]. China [Rod18, Rod17, Set17a]. Claudia [Ano14d]. clubhouse [Dav12]. co [TY13b]. co-director [TY13b]. code [Doy16a, PS13, Edi14]. cohort [GO19, SC18]. College [Ano15a, Ham15, Kau15, Shu15]. coming [PS13]. commitment [Mor19b]. committees [She17]. Community [Ano15a, Wal12a]. CompEd [ZCLR19]. computational [Wal13]. Computer [AO12, C^{+}11, Tym14a, Wol11, Bar17a, Far12, Tym14c, Tym15, RF11, Ric12, SR13]. Computing [SM17, ZCLR19, KD14, LB18, RS16, Tym14b, Wal13, Guz11]. Conference [ZCLR19, BT15, CCS13, Cut14, DCS15a, IZW18, MJ18, Dag15, DE15a, Set17a, SM17]. conferences [Wal11b]. Confessions [Rie11]. contribute [CSD14]. Coordinating [Guz11]. copyrights [Wal11a]. Cornell [Kau15]. Council [Guz11, Fit15a, Owe11a, Owe12a]. County [Ano15a]. CRA [GO19, RHM18, SC18]. CRA-W [GO19, RHM18, SC18]. CS [AB11b, Bar12, Bar17c, Edi14, Joy11]. CSE [RFW11]. CSED [Chu12, Far12]. CSEdWeek [AS13b]. CSforAll [Mor19b]. CSTA [Ste14]. curricula [SR13]. cyber [Fit15b].

Dallas [C^{+}11]. Daniels [Fit17a]. dates [SIG18a, SIG18b, SIG19a, SIG19b]. days [Rod14]. Deadlines [SIG18a, SIG18b, SIG19a, SIG19b]. decade [PQ19]. department [DeL14a, DeL14b, Mah18]. Development [Max17]. dialog [Wal13]. did [AS13b, DN14, Doy16a]. Digital [Owe11b]. director [TY13b]. discrimination [Set18a]. diverse [PS16]. diversity [KB15]. doctoral [RS16]. Drexel [Bol15].

Ed [Bar12, Edi14, Joy11, RF11, Owe11a]. Education [C^{+}11, Fit15a, Guz11, Owe12a, Ric12, SM17, Tym14a, ZCLR19, Bar17a, Fit15b, RS16, Tym14b, Tym15, Wol11]. educator [DD17, Tym14c]. educators [DH19, RD15]. election [McC16, Rod19a]. elections [MR15]. Elegit [Mus17]. encouraged [Wal13]. engineering [Max17]. enrollments [KB15]. equity [Bar17c]. error [TY13b]. European [Jul13]. experience [DN14].

faculty [PS16]. First [ZCLR19]. Florida [Cha15]. Force [Had15]. fuels [Far12]. fund [Set11, Set15]. funding [Tym14b]. future [Far12].

Gail [Ano17b]. game [Max17]. Get [Fit14]. gets [Bar15]. getting [SR13]. GHC [Dav12]. Glasgow [CSD14]. Global [Rod18, ZCLR19, IZW18, Joh10]. Grace [Ada11]. grad [GO19, SC18]. Grant [BF14, Dec18b, Dec19b, RS13, Set14a, Set16a]. grants [Bal11, Bal12, Dec17, Dec18a, Fit15c, Fit16, Fit17b, Fit18b, Fit18a, Set14b, SS13, She15a, She18, She19b, She19a]. group [AB11a, Rie11]. groups [Bar13]. growth [RS16]. guide [Tym14c].

Haifa [Wal12a]. harassment [Set18a]. Help [DEA18]. Highlight [Por18]. Highlights [BPG15, BP16]. history [Doy15]. Hopper [Ada11]. Hour [Doy16a, PS13, Edi14]. HS [ACG^{+}13].

ICER [CT17, CSS12, CCS13, Cut13, CDS13, Cut14, CSD14, DCS15a, DCS15b, MKMP18, MMK18, MP19a, MPRM19, MP19b, SDCT16a, SDCT16c, SDCT16b, She16a, TCSM17, TCM^{+}17]. ICER2015 [DSC15]. improved [Cas16]. improvement [Max17]. information [TY13b]. initiatives [Dec19c]. Inroads [IS13]. interface [Owe11b]. International [Cha15, SM17]. interview [LB18]. involved [Fit14, SR13]. IRB [Jad16]. iSIGCSE [Dix15]. Israel [Wal12a]. ITiCSE [AB11a, Bar13, CDCP14, Car13, Cle16, CCV16, CCV17, Dag15, DJ15, DG17, Gol17, Gol18, LGE12, MS19, Rod14, Swe13, Wal12a].

J [Mil19]. Jane [Dav12]. JavaScript [Wal12c]. JENUI [Jul13]. join [BT15, DCS15b]. junkie [Rie11].

K-12 [Wol11]. kick [Chu12]. kick-off [Chu12]. Kid [FP12]. Ko [Mil19]. Koli [FI19, JI18, MJ18, SM17]. languages [PQ19]. later [Dav12]. launches [IZW18]. Lawrence [Chu12]. leadership [Sob14]. Learn [PB16]. length [HZPQH18a]. leverage [DN14]. Library [Owe11b]. Life [TY13b]. like [AB11a]. lists [Mor19a, TY13a]. look [CDS13].

mailing [Mor19a]. March [C^{+}11]. Margolis [Dav12]. Mathematics [Wal13]. Mats [Fit17a]. May [Rod18, She19a]. meeting [PS16, She14, She15b, She16b]. meets [Owe12a]. Member [AS13a, Ano13a, Ano13b, Ano14c, Ano14d, Ano14a, Ano14b, Ano15b, Ano15c, Ano15d, Ano15e, Ano17a, Ano17b, Dav18, Dav19a, Dav19b, Doy16b, Doy16c, Doy17a, Doy17b, Doy18, Edi13, Fit17a, Kau16a, Kau16b, Mil19, Por18, PG19, Swe12]. members [Ste14]. mentions [PQ19]. messages [TY13b]. Multiple [PQH17].

NCWIT [Alv12, Bra18]. need [PS16]. needs [KD14]. networking [Wal12a]. news [Wol11]. next [Rod19b]. nomination [Tym14c]. November [BF14, Fit18b]. NSF [Tym14b]. numbers [Wal11c, Wal11d, Wal12b, Wal12c]. Nurturing [Sob14].

odds [KB15]. off [Chu12]. Omaha [DCS15b]. Open [Dec19a]. opportunities [Bar17b]. organization [BT15]. our [BT15].

Paper [Wal11b, Alv13, HZPQH18a, HZPQH18b, PQH17, Wal10]. papers [DW19, Dec19d]. part [Ano17b, Fit17a]. participate [AB11a]. participation [BT15, BPG15, Bar15, BP16, Bar17b, Cut19, MP19a, PS16, PB16, PT18, RHM18, Tow16]. partnership [ACG^{+}13]. past [Dec19a]. perspective [Alv12]. plagiarism [Wal11a]. plans [Tow16]. policy [Set17b, Set18a]. portfolio [AO12]. portfolio-based [AO12]. preparing [KD14]. Preview [Cut13, AT16a, BG18a, CDCP14, Cam13, CCV16, DJ15, DG17, DE15b, DSC15, DCS15b, EC17, Gol18, HPQ19, JI18, MKMP18, MPRM19, MS19, ND14, Set18b, SDCT16a, SDCT16c, SDCT16b, TCSM17, TCM^{+}17]. principles [AO12, AB11b]. proceedings [C^{+}11, Dec19a]. process [Cas16, Max17]. professional [BT15]. program [Dec18b, Dec19b, RS13, Set11, Set14a, Set16a]. programming [PQ19]. project [AB11b, Fit15b, Fit16, Fit18a, SS13, She15a, She18]. Projects [BF14, Bal11, Bal12, Bal13, Fit15c, Fit17b, Fit18b, She19b, She19a]. puzzle [Owe11a].

ready [SR13]. recap [Car13, DN14]. reform [ACG^{+}13]. Report [ZCLR19, BG18b, CCS13, CCV17, Cut14, Dag15, DE15a, DCS15a, Fit15a, HPQHZ19, KD14, LGE12, McC12, MJ18, Owe11a, Rod15b, Rod16, Set17c, Set18b, She14, She15b, She16b]. required [Jad16]. research [BPG15, BP16, PB16, PT18, RS16]. resources [Ste14]. RESPECT [BT15, Bar15, PSTT17, PT18, Tow16]. RESPECT’15 [BPG15]. RESPECT’16 [BP16, PB16]. responsibility [Bar17c]. review [AT16b, CT17, Gol17, MMK18, MP19b, She16a, SM17]. reviews [Jad16]. Revised [Set17b]. Robofest [Chu12]. round [She19a]. Roundtable [Mah16, DH19, Mah18].

scale [IZW18]. Science [C^{+}11, RF11, Ric12, Tym14a, AO12, Bar17a, Far12, Tym14c, Tym15, Wal13, SR13, Wol11]. Scotland [CSD14]. shines [Ric12]. showcase [Bar12]. SIGCE [PQ19]. SIGCSE [Tym14a, AT16a, AT16b, Alv12, Alv13, Bal11, Bal13, Bar15, BG18a, BG18b, Cam13, Cas16, DD17, DeL14a, DeL14b, DE15a, DE15b, Dec17, DEA18, Dec18a, DWS19, Dec19a, Dec19c, Dec19b, DW19, Dec19d, DH19, DN14, Doy15, Doy16d, EC17, Fit14, Fit15c, Fit16, Fit17b, Fit18b, Fit18a, FP12, HPQ19, HPQHZ19, HZPQH18a, HZPQH18b, IZW18, Joh10, KB15, Mah16, Mah18, McC12, MR15, McC16, MS18, Mor19b, Mor19a, ND14, PQH17, RD15, RS13, Rod15a, Rod15b, Rod16, Rod17, Rod18, Rod19a, Rod19b, Set11, Set14b, Set16b, Set17c, Set18b, SFM19, SS13, She14, She15b, She15a, She16b, She17, She18, She19a, Sim13, Ste14, TY13a, TY13b, Tym14c, Tym15, Wal11b, Wal11c, Wal11d, Wal12b, Wal12c, ZSH^{+}19]. SIGCSE’11 [C^{+}11]. skills [Sob14]. slate [Rod19a]. software [Max17]. Southern [Ano15f]. speaker [Set11]. Special [Bal12, BF14, She19b, Bal11, Bal13, Fit15c, Fit16, Fit17b, Fit18b, Fit18a, SS13, She15a, She18, She19a]. Spotlight [Dav18, AS13a, Ano13a, Ano13b, Ano14c, Ano14d, Ano14a, Ano14b, Ano15b, Ano15c, Ano15d, Ano15e, Ano17a, Ano17b, Dav12, Dav19a, Dav19b, Doy16b, Doy16c, Doy17a, Doy17b, Doy18, Edi13, Fit17a, Kau16a, Kau16b, Mil19, PG19, Swe12]. Still [CSD14]. student [Sob14]. students [KD14]. submissions [FI19, Wal11b, ZSH^{+}19]. summary [Mus17]. summer [Tow16]. survey [Joy11]. Symposium [AT16a, Alv13, BG18b, Cam13, C^{+}11, EC17, FP12, She15b, AT16b, BG18a, Cas16, DE15b, DW19, Dec19d, Doy15, HPQHZ19, HZPQH18b, PQ19, ZSH^{+}19]. symposiums [DEA18]. Szabo [Ano14d].

Tech [Chu12]. Technical [C^{+}11, AT16b, Cas16, DEA18, HPQHZ19, PQ19, ZSH^{+}19]. ten [DW19, Dec19d]. test [DWS19]. Texas [C^{+}11]. Three [Rod14]. time [CSD14, DWS19, DW19, Dec19d]. top [DW19, Dec19d]. Training [DeL14a, DeL14b]. Travel [Set15, Dec17, Dec18a, Dec18b, Dec19b, RS13, Set14b, Set14a, Set16a]. Turing [Owe12b, Set17a]. types [PQH17].

U.S. [Had15, KD14]. UK [Swe13]. UNC [Bar12]. University [Ano15f, Bol15, Cha15]. unlocking [Dav12]. upcoming [SIG18a, SIG18b, SIG19a, SIG19b]. Update [AB11b]. USA [C^{+}11]. Utah [Ano15f]. view [Jul13]. vision [Joh10]. Volunteering [SFM19]. volunteers [Rod15a]. Vote [McC16, Rod19b].

W [GO19, RHM18, SC18]. wants [RFW11]. website [IS13]. wednesday [DH19]. Week [Edi14, Chu12, Far12, RFW11, Bar12, Joy11, RF11, Ric12]. welcome [Set16b]. winners [Dec19d, Doy16d]. WiPSCE [Cut19]. women [GO19]. workforce [KD14]. working [AB11a, Bar13, Rie11]. workshop [DD17, RD15].

year [McC16]. years [Dav12].

References

Adams:2011:WLP

[AB11a] Liz Adams and Jürgen Börstler. What it’s like to participate in an ITiCSE working group. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 43(1):6, March 2011. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:2011:UCP

[AB11b] Owen Astrachan and Amy Briggs. Update on the CS Principles project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 43(4):2, October 2011. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:2013:NPR

[ACG^{+}13] Owen Astrachan, Gail Chapman, Joanna Goode, Brook Osborne, and Pat Yongpradit. A new partnership to reform HS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(4):4, October 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:2011:CGH

[Ada11] Joel Adams. Calvin’s Grace Hopper birthday celebration. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 43(1):2, March 2011. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Alvarado:2012:NSP

[Alv12] Christine Alvarado. NCWIT, from a SIGCSE perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(3):9, July 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Alvarado:2013:SSB

[Alv13] Christine Alvarado. SIGCSE Symposium 2013: best paper award. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(2):9, April 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:2013:MSa

[Ano13a] Anonymous. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(2):3–5, April 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:2013:MSb

[Ano13b] Anonymous. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(3):4–5, July 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:2014:MSa

[Ano14a] Anonymous. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(1):5–7, January 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:2014:MSb

[Ano14b] Anonymous. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(3):4–6, July 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:2014:MS

[Ano14c] Anonymous. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(4):5–7, October 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:2014:MSC

[Ano14d] Anonymous. Member spotlight: Claudia Szabo. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(2):4–5, April 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:2015:BCC

[Ano15a] Anonymous. Bucks County Community College. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(1):4, January 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:2015:MSa

[Ano15b] Anonymous. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(1):8–10, January 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:2015:MSb

[Ano15c] Anonymous. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(2):3–5, April 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:2015:MSc

[Ano15d] Anonymous. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(3):4–6, July 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:2015:MSd

[Ano15e] Anonymous. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(4):5, October 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:2015:SUU

[Ano15f] Anonymous. Southern Utah University. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(1):4, January 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:2017:MS

[Ano17a] Anonymous. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(3):4–6, July 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:2017:MSP

[Ano17b] Anonymous. Member spotlight part 1: Gail Chapman. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(2):5–11, April 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:2012:CSP

[AO12] Owen Astrachan and R. Brook Osborne. Computer science principles: portfolio-based assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(4):4, October 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Alvarado:2013:MS

[AS13a] Christine Alvarado and Z. Sweedyk. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(1):5–7, January 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Alvarado:2013:WWD

[AS13b] Christine Alvarado and Z. Sweedyk. What we did: CSEdWeek 2012. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(1):2–4, January 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Alphonce:2016:SSP

[AT16a] Carl Alphonce and Jodi Tims. SIGCSE Symposium 2016 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(1):5, January 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Alphonce:2016:STS

[AT16b] Carl Alphonce and Jodi Tims. SIGCSE technical symposium 2016 review. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(3):8, July 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baldwin:2011:SSP

[Bal11] Doug Baldwin. SIGCSE special projects grants awarded. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 43(4):5, October 2011. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baldwin:2012:SPG

[Bal12] Doug Baldwin. Special projects grants awarded. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(3):6, July 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baldwin:2013:SSP

[Bal13] Doug Baldwin. SIGCSE special projects awards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(1):7, January 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnes:2012:CEW

[Bar12] Tiffany Barnes. CS Ed Week showcase at UNC Charlotte. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(1):7, January 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnes:2013:IWG

[Bar13] David J. Barnes. ITiCSE 2013: working groups. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(3):8, July 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnes:2015:SBB

[Bar15] Tiffany Barnes. SIGCSE BP: broadening participation gets RE-SPECT. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(3):8, July 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnes:2017:BAC

[Bar17a] Tiffany Barnes. Broadening access to computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(4):11–12, October 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnes:2017:BPO

[Bar17b] Tiffany Barnes. Broadening participation opportunities in 2017. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(1):10–11, January 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnes:2017:CAE

[Bar17c] Tiffany Barnes. CS for all, equity, and responsibility. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(2):18, April 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnes:2014:NSP

[BF14] Tiffany Barnes and Sue Fitzgerald. November 2013 Special Projects Grant Awards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(1):9, January 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnes:2018:SSP

[BG18a] Tiffany Barnes and Dan Garcia. SIGCSE symposium 2018 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(1):3–6, January 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnes:2018:SSR

[BG18b] Tiffany Barnes and Dan Garcia. SIGCSE Symposium 2018 report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(2):2–3, April 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boland:2015:DU

[Bol15] Kerry Boland. Drexel University. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(1):4, January 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnes:2016:HBP

[BP16] Tiffany Barnes and Jamie Patent. Highlights of broadening participation research at RESPECT’16. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(4):5, October 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnes:2015:HBP

[BPG15] Tiffany Barnes, Jamie Payton, and Mark Guzdial. Highlights of broadening participation research at RESPECT’15. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(4):3, October 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bradshaw:2018:N

[Bra18] Eva Bradshaw. NCWIT. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(1):12, January 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnes:2015:BPR

[BT15] Tiffany Barnes and George Thiruvathukal. Broadening participation and RESPECT: join our new professional organization and conference! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(1):7, January 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cortina:2011:SPA

[C^{+}11] Thomas J. Cortina et al., editors. SIGCSE’11: proceedings of the 42nd ACM Technical Symposium on Computer Science Education: March 9–12, 2011, Dallas, Texas, USA, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2011. CODEN SIGSD3. ISBN 1-4503-0500-8. ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN ????

Camp:2013:SSP

[Cam13] Tracy Camp. SIGCSE Symposium 2013 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(1):8, January 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carter:2013:IR

[Car13] Janet Carter. ITiCSE 2013: recap. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(3): 7, July 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Casperson:2016:IPS

[Cas16] Michael Casperson. An improved process for the SIGCSE technical symposium. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(2):5, April 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:2013:ICR

[CCS13] Alison Clear, Quintin Cutts, and Beth Simon. ICER 2013 conference report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(4):7, October 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:2016:IP

[CCV16] Alison Clear and Ernesto Cuadros-Vargas. ITiCSE 2016 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(1):6, January 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:2017:IR

[CCV17] Alison Clear and Ernesto Cuadros-Vargas. ITiCSE 2016 report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(1):4, January 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cajander:2014:IP

[CDCP14] Åsa Cajander, Mats Daniels, Tony Clear, and Arnold Pears. ITiCSE preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(2):8–9, April 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cutts:2013:LAI

[CDS13] Quintin Cutts, Brian Dorn, and Beth Simon. A look ahead to ICER 2014. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(4):8, October 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Charters:2015:FIU

[Cha15] Maria Charters. Florida International University. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(1):3, January 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chung:2012:RKC

[Chu12] Chan Jin Chung. Robofest kick-off for CSED week at Lawrence Tech. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(1):6, January 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:2016:I

[Cle16] Alison Clear. ITiCSE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(2):4, April 2016. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cutts:2014:STC

[CSD14] Quintin Cutts, Beth Simon, and Brian Dorn. Still time to contribute to ICER 2104, Glasgow, Scotland. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(2):9, April 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:2012:I

[CSS12] Alison Clear, Kate Sanders, and Beth Simon. ICER 2012. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(3):4, July 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chinn:2017:IR

[CT17] Donald Chinn and Josh Tenenberg. ICER review. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(4):4–5, October 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cutts:2013:IP

[Cut13] Quintin Cutts. ICER 2013 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(3):9, July 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cutts:2014:ICR

[Cut14] Quintin Cutts. ICER conference report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(4):4, October 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cutts:2019:WCP

[Cut19] Quintin Cutts. WiPSCE 2019 call for participation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(2):8, April 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dagiene:2015:ICR

[Dag15] Valentina Dagiene. ITiCSE Conference report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(3):2–3, July 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davis:2012:GSJ

[Dav12] Janet Davis. GHC spotlight: Jane Margolis, unlocking the clubhouse 10 years later. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(4):3, October 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davis:2018:MS

[Dav18] Karen Davis. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(4):6–8, October 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davis:2019:MSa

[Dav19a] Karen Davis. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(2):9–10, April 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davis:2019:MSb

[Dav19b] Karen Davis. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(3):8–11, July 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dorn:2015:ICR

[DCS15a] Brian Dorn, Quintin Cutts, and Judy Sheard. ICER conference report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(4):2–3, October 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dorn:2015:IPJ

[DCS15b] Brian Dorn, Quintin Cutts, and Judy Sheard. ICER preview: join us in Omaha. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(3):3–4, July 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Danyluk:2017:SNE

[DD17] Andrea Danyluk and Zachary Dodds. SIGCSE 2018 new educator workshop. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(4):6, October 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Decker:2015:SCR

[DE15a] Adrienne Decker and Kurt Eiselt. SIGCSE Conference report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(3):2, July 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Decker:2015:SSP

[DE15b] Adrienne Decker and Kurt Eiselt. SIGCSE symposium preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(1):4–5, January 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Decker:2018:HUC

[DEA18] Adrienne Decker, Kurt Eiselt, and Carl Alphonce. Help us celebrate 50 SIGCSE technical symposiums. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(3):2, July 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Decker:2017:STG

[Dec17] Adrienne Decker. 2017 SIGCSE travel grants. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(4):7, October 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Decker:2018:STG

[Dec18a] Adrienne Decker. 2018 SIGCSE travel grants. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(3):11, July 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Decker:2018:TGP

[Dec18b] Adrienne Decker. 2018 travel grant program awards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(1):6–8, January 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Decker:2019:OAP

[Dec19a] Adrienne Decker. Open access to past SIGCSE proceedings. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(1):10, January 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Decker:2019:STG

[Dec19b] Adrienne Decker. SIGCSE 2018 travel grant program awards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(2):12–13, April 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Decker:2019:SBI

[Dec19c] Adrienne Decker. SIGCSE board initiatives. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(4):2–3, November 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3371394.3371395.

Decker:2019:STTb

[Dec19d] Adrienne Decker. SIGCSE top ten symposium papers of all time award winners. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(2):14–15, April 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

DeLoatch:2014:TDCa

[DeL14a] Sandra J. DeLoatch. Training for department chairs at SIGCSE 2015. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(2):6, April 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

DeLoatch:2014:TDCb

[DeL14b] Sandra J. DeLoatch. Training for department chairs at SIGCSE 2015. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(4):8, October 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davoli:2017:IP

[DG17] Renzo Davoli and Michael Goldweber. ITiCSE 2017 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(1):4–5, January 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dodds:2019:NEW

[DH19] Zachary Dodds and Diane Horton. New educators wednesday roundtable at SIGCSE 2020. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(4):3, November 2019. CODEN SIGSD3. ISSN 0097-8418 (print),

2331-3927 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3371394.3371396.

Dixit:2015:I

[Dix15] Arati Dixit. iSIGCSE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(2):7, April 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dagiene:2015:IP

[DJ15] Valentina Dagiene and Tatjana Jevsikova. ITiCSE preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(2):2, April 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dougherty:2014:SRD

[DN14] John P. Dougherty and Kris Nagel. SIGCSE 2014 recap: did you experience leverage at SIGCSE? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(2): 2–3, April 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Doyle:2015:SSH

[Doy15] Maureen Doyle. SIGCSE symposium history. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(4):7–8, October 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Doyle:2016:HCW

[Doy16a] Maureen Doyle. Hour of code: what you did. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(1):2–3, January 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Doyle:2016:MSa

[Doy16b] Maureen Doyle. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(2):6–7, April 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Doyle:2016:MSb

[Doy16c] Maureen Doyle. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(4):6–7, October 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Doyle:2016:SAW

[Doy16d] Maureen Doyle. SIGCSE award winners. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(2):2, April 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Doyle:2017:MSa

[Doy17a] Maureen Doyle. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(1):7–8, January 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Doyle:2017:MSb

[Doy17b] Maureen Doyle. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(4):7–11, October 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Doyle:2018:MS

[Doy18] Maureen Doyle. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(2):4–8, April 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dorn:2015:IP

[DSC15] Brian Dorn, Judy Sheard, and Quintin Cutts. ICER2015 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(1):6, January 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Decker:2019:STTa

[DW19] Adrienne Decker and Mark Allen Weiss. SIGCSE top ten symposium papers of all time award. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(1):9, January 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Decker:2019:AST

[DWS19] Adrienne Decker, Mark Allen Weiss, and Judithe Sheard. Announcing the SIGCSE test of time award. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(2):17, April 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Edwards:2017:SSP

[EC17] Stephen H. Edwards and Michael E. Caspersen. SIGCSE Symposium 2017 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(1):3, January 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Editors:2013:MS

[Edi13] Editors. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(4):2–3, October 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Editors:2014:CEW

[Edi14] Editors. CS Ed Week 2013: the Hour of Code. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(1):2–4, January 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Farmer:2012:CWC

[Far12] Ruthe Farmer. CSED week: computer science fuels the future. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(4):5, October 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Falkner:2019:KCC

[FI19] Nick Falkner and Petri Ihantola. Koli Calling 2019 call for submissions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(3):3–4, July 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fitzgerald:2014:GIS

[Fit14] Sue Fitzgerald. Get involved with SIGCSE! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(3):10, July 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fitzgerald:2015:AEC

[Fit15a] Sue Fitzgerald. ACM Education Council report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(1):7, January 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fitzgerald:2015:CEP

[Fit15b] Sue Fitzgerald. The cyber education project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(2):8–9, April 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fitzgerald:2015:SSP

[Fit15c] Sue Fitzgerald. SIGCSE special projects grants. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(2):6, April 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fitzgerald:2016:SSP

[Fit16] Sue Fitzgerald. SIGCSE special project grants. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(3):3–4, July 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fitzgerald:2017:MSP

[Fit17a] Sue Fitzgerald. Member spotlight part 2: Mats Daniels. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(2):11–14, April 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fitzgerald:2017:SSP

[Fit17b] Sue Fitzgerald. SIGCSE special projects grants. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(1):9–10, January 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fitzgerald:2018:SSP

[Fit18a] Sue Fitzgerald. 2018 SIGCSE special project grants. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(3):12, July 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fitzgerald:2018:NSS

[Fit18b] Sue Fitzgerald. November 2017 SIGCSE special projects grants. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(1):11–12, January 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fox:2012:SSK

[FP12] Susan Fox and Sarah Monisha Pulimood. SIGCSE Symposium Kids’ Camp. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(3):5, July 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gorson:2019:CWG

[GO19] Jamie Gorson and Alannah Oleson. CRA-W grad cohort for women. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(3):7–8, July 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldweber:2017:IR

[Gol17] Michael Goldweber. ITiCSE review. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(4): 3–4, October 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldweber:2018:IP

[Gol18] Michael Goldweber. ITiCSE 2018 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(2):8–9, April 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Guzdial:2011:CEC

[Guz11] Mark Guzdial. Computing Education Coordinating Council (CECC). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 43(1):7, March 2011. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hadfield:2015:UAF

[Had15] Steve Hadfield. U.S. Air Force Academy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(1):3, January 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamid:2015:BC

[Ham15] Nadeem Abdul Hamid. Berry College. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(1):2, January 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hawthorne:2019:SP

[HPQ19] Elizabeth K. Hawthorne and Manuel A. Pérez-Quiñones. SIGCSE 2019 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(1):2, January 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hawthorne:2019:STS

[HPQHZ19] Elizabeth K. Hawthorne, Manuel A. Pérez-Quiñones, Sarah Heckman, and Jian Zhang. SIGCSE technical symposium 2019 report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(2):2–4, April 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Heckman:2018:SPL

[HZPQH18a] Sarah Heckman, Jian Zhang, Manuel A. Peérez-Quiñones, and Elizabeth K. Hawthorne. SIGCSE 2019 paper length change. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(2):4, April 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Heckman:2018:WSS

[HZPQH18b] Sarah Heckman, Jian Zhang, Manuel A. Pérez-Quiñones, and Elizabeth K. Hawthorne. What is a SIGCSE symposium paper? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(3):3, July 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:2013:AIW

[IS13] John Impagliazzo and Amber Settle. ACM Inroads: the website. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(1):10, January 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:2018:SLN

[IZW18] John Impagliazzo, Ming Zhang, and Xi Wu. SIGCSE launches new conference on a global scale. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(4):2–3, October 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jadud:2016:IRR

[Jad16] Matthew Jadud. IRB reviews required. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(3):9, July 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Joy:2018:KCP

[JI18] Mike Joy and Petri Ihantola. Koli Calling 2018 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(4):3–4, October 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Johnson:2010:SGV

[Joh10] Julie Johnson. SIGCSE: global vision. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 42(3): 2, September 2010. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Joyce:2011:CEW

[Joy11] Dan Joyce. CS Ed Week activities survey. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 43(1):2, March 2011. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Julia:2013:EVJ

[Jul13] José Juan Miró Julià. The European view: JENUI. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(3):6, July 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kauchak:2015:CC

[Kau15] David Kauchak. Cornell College. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(1):2, January 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kauchak:2016:MSa

[Kau16a] David Kauchak. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(1):8–9, January 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kauchak:2016:MSb

[Kau16b] David Kauchak. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(3):6–7, July 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kaczmarczyk:2015:SBE

[KB15] Lisa Kaczmarczyk and Tiffany Barnes. SIGCSE BP: enrollments and diversity at odds? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(2):8, April 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kaczmarczyk:2014:ARP

[KD14] Lisa Kaczmarczyk and Renee Dopplick. ACM report: preparing students for computing workforce needs in the U.S. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(2):8, April 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ladner:2018:IAC

[LB18] Richard Ladner and Tiffany Barnes. An interview with access computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(1):9–11, January 2018. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lapidot:2012:ICR

[LGE12] Tami Lapidot and Judith Gal-Ezer. ITiCSE 2012: chairs’ report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(3):2, July 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Maher:2016:SCR

[Mah16] Mary Lou Maher. SIGCSE 2017 Chairs roundtable. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(4):7–8, October 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Maher:2018:SDC

[Mah18] Mary Lou Maher. SIGCSE 2019 department chairs roundtable. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(4):11, October 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Maxim:2017:DSE

[Max17] Bruce Maxim. Development of a software engineering process improvement game. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(2):16–17, April 2017. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:2012:SCA

[McC12] Renee McCauley. SIGCSE chair’s annual report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(3):6–7, July 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:2016:VEY

[McC16] Renée McCauley. Vote!: 2016 is election year for SIGCSE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(2):2, April 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Miller:2019:MSA

[Mil19] Jeffrey Miller. Member spotlight: Amy J. Ko. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(4):6–8, November 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3371394.3371399.

Montero:2018:KCC

[MJ18] Calkin Suero Montero and Mike Joy. Koli Calling 2017 conference report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(2):10–11, April 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Malmi:2018:IP

[MKMP18] Lauri Malmi, Ari Korhonen, Robert McCartney, and Andrew Petersen. ICER 2018 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(2):9–10, April 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCartney:2018:IR

[MMK18] Robert McCartney, Lauri Malmi, and Ari Korhonen. ICER 2018 review. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(4):4–5, October 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Morrison:2019:SML

[Mor19a] Briana Morrison. SIGCSE mailing lists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(3): 11–12, July 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Morrison:2019:SCC

[Mor19b] Briana B. Morrison. SIGCSE’s commitment to CSforAll. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(1):9–10, January 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCartney:2019:ICP

[MP19a] Robert McCartney and Andrew Petersen. ICER 2019 call for participation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(1):3–4, January 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCartney:2019:IR

[MP19b] Robert McCartney and Andrew Petersen. ICER 2019 review. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(4):5–6, November 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3371394.3371398.

McCartney:2019:IP

[MPRM19] Robert McCartney, Andrew Petersen, Anthony Robins, and Adon Moskal. ICER 2019 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(2):7, April 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:2015:SE

[MR15] Renée McCauley and Susan Rodger. 2016 SIGCSE elections. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(4):4, October 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Morrison:2018:CSA

[MS18] Briana Morrison and Amber Settle. Celebrating SIGCSE’s 50th anniversary! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(1):2–3, January 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McDermott:2019:IP

[MS19] Roger McDermott and Bruce Scharlau. ITiCSE 2019 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(2):6, April 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Musicant:2017:ES

[Mus17] David Musicant. Elegit summary. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(2):17, April 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nagel:2014:SP

[ND14] Kris Nagel and J. D. Dougherty. SIGCSE 2014 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(1):9–11, January 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Owens:2011:ECE

[Owe11a] Barabara Boucher Owens. Ed Council and Ed Board: the puzzle and the report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 43(4):4, October 2011. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Owens:2011:NAD

[Owe11b] Barbara Boucher Owens. The NEW ACM Digital Library interface. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 43(2):6, June 2011. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Owens:2012:AEC

[Owe12a] Barbara Boucher Owens. ACM Education Council meets. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(3):7–8, July 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Owens:2012:ATC

[Owe12b] Barbara Boucher Owens. ACM Turing Centenary celebration. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(3):8, July 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Payton:2016:LAB

[PB16] Jamie Payton and Tiffany Barnes. Learn about broadening participation: research at RESPECT’16. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(3):5, July 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Porter:2019:MS

[PG19] Leo Porter and Mark Guzdial. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(1):4–7, January 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Porter:2018:MH

[Por18] Leo Porter. Member highlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(3):4–8, July 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Perez-Quinones:2019:STS

[PQ19] Manuel A. Pérez-Quiñones. SIGCE technical symposium programming languages mentions by decade. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(2):4–5, April 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Perez-Quinones:2017:MPT

[PQH17] Manuel A. Perez-Quinones and Elizabeth K. Hawthorne. Multiple paper types for SIGCSE 2018. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(3):7, July 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Partovi:2013:HCC

[PS13] Hadi Partovi and Mehran Sahami. The hour of code is coming! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(4):5, October 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Payton:2016:BPM

[PS16] Jamie Payton and Richard Souvenir. Broadening participation: meeting the need for diverse faculty. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(1): 10, January 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Payton:2017:R

[PSTT17] Jamie Payton, Felesia Stukes, George K. Thiruvathul, and Jakita O. Thomas. RESPECT ’18. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(3): 6, July 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Payton:2018:ABP

[PT18] Jamie Payton and George K. Thiruvathukal. Advancing broadening participation research at RESPECT. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(3):9– 10, July 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reed:2015:SNE

[RD15] David Reed and Andrea Danyluk. SIGCSE 2016 new educators workshop. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(4):6, October 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Richardson:2011:WCS

[RF11] Debra Richardson and Ruthe Farmer. Why Computer Science Ed Week? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 43(1):3, March 2011. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Richardson:2011:CWW

[RFW11] Debra Richardson, Ruthe Farmer, and Cameron Wilson. CSE week wants you. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 43(3):6, September 2011. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Russell:2018:BPC

[RHM18] Erik Russell, Julia Hirschberg, and Margaret Martonosi. Broadening participation: CRA-W. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(2):12, April 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Richardson:2012:CSE

[Ric12] Debra Richardson. Computer Science Education Week shines in 2011. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(3):10, July 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Riedesel:2011:CWG

[Rie11] Charles Riedesel. Confessions of a working group junkie. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 43(1):6, March 2011. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:2014:TDI

[Rod14] Susan Rodger. Three days at ITiCSE 2014. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(3):6–8, July 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:2015:NSV

[Rod15a] Susan Rodger. New SIGCSE volunteers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(2):5, April 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:2015:SCR

[Rod15b] Susan Rodger. SIGCSE Chair’s report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(3): 6–7, July 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:2016:SCR

[Rod16] Susan Rodger. SIGCSE Chair report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(3): 9–10, July 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:2017:NSC

[Rod17] Susan Rodger. New SIGCSE chapter: China. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(1):2, January 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:2018:SGC

[Rod18] Susan Rodger. SIGCSE global — Chengdu, China, May 17–19, 2019. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(3):8–9, July 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:2019:SBE

[Rod19a] Susan Rodger. SIGCSE Board 2019–2022 election slate. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(1):8, January 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:2019:VNS

[Rod19b] Susan Rodger. Vote for the next SIGCSE board. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(2):11, April 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:2013:STG

[RS13] Susan Rodger and Amber Settle. SIGCSE travel grant program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(3):6, July 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robins:2016:GCE

[RS16] Anthony Robins and R. Benjamin Shapiro. The growth of computing education doctoral research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(4):3, October 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Salac:2018:CWG

[SC18] Jean Paula Salac and Jieun Chon. CRA-W grad cohort. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(4):9–10, October 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2016:IP

[SDCT16a] Judy Sheard, Brian Dorn, Donald Chinn, and Josh Tenenberg. ICER 2016 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(3):4, July 2016. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2016:IPb

[SDCT16b] Judy Sheard, Brian Dorn, Donald Chinn, and Josh Tenenberg. ICER preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(2):3, April 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2016:IPa

[SDCT16c] Judy Sheard, Brian Dorn, Donald Chinn, and Josh Tenenburg. ICER 2016 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(1):7, January 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Settle:2011:SSF

[Set11] Amber Settle. SIGCSE speaker’s fund program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 43(4):5, October 2011. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Settle:2014:TGP

[Set14a] Amber Settle. 2014 travel grant program awards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(2):5–6, April 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Settle:2014:STG

[Set14b] Amber Settle. SIGCSE travel grants. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(4):7, October 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Settle:2015:TFA

[Set15] Amber Settle. Travel fund awardees. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(3): 9–10, July 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Settle:2016:TGP

[Set16a] Amber Settle. 2016 travel grant program awards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(1):3–4, January 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Settle:2016:SBC

[Set16b] Amber Settle. SIGCSE Board Chair welcome. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(4):2, October 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Settle:2017:ATC

[Set17a] Amber Settle. ACM Turing 50th Celebration Conference in China. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(3):2–3, July 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Settle:2017:RAA

[Set17b] Amber Settle. Revised ACM authorship policy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(1):9, January 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Settle:2017:SCR

[Set17c] Amber Settle. SIGCSE chair report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(3): 3–4, July 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Settle:2018:APA

[Set18a] Amber Settle. ACM policy against discrimination and harassment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(1):9, January 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Settle:2018:SCR

[Set18b] Amber Settle. SIGCSE chair’s report preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(3):10–11, July 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Settle:2019:VS

[SFM19] Amber Settle, Sue Fitzgerald, and Briana Morrison. Volunteering with SIGCSE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(2):16–17, April 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2014:SBM

[She14] Judy Sheard. SIGCSE business meeting report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(2):3, April 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2015:SSP

[She15a] Judy Sheard. SIGCSE special project grants awards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(3):8–9, July 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2015:SSB

[She15b] Judy Sheard. SIGCSE Symposium 2015: business meeting report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(2):9, April 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2016:IR

[She16a] Judy Sheard. ICER 2016 review. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(4):4–5, October 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2016:SBM

[She16b] Judy Sheard. SIGCSE business meeting report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(2):8, April 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2017:SC

[She17] Judith Sheard. SIGCSE committees. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(4): 12, October 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2018:SSP

[She18] Judy Sheard. 2018 SIGCSE special project grants. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(4):12, October 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2019:SSP

[She19a] Judith Sheard. SIGCSE special projects grants awards: May 2019 round. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(4):4–5, November 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3371394.3371397.

Sheard:2019:SPG

[She19b] Judithe Sheard. Special projects grants. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(2): 15–16, April 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shuai:2015:BC

[Shu15] Xiangyin Shuai. Bard College. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(1):2–3, January 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Staff:2018:DUDa

[SIG18a] SIGCSE Staff. Deadlines and upcoming dates! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(3):13, July 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Staff:2018:DUDb

[SIG18b] SIGCSE Staff. Deadlines and upcoming dates! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 50(4):12, October 2018. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Staff:2019:DUDa

[SIG19a] SIGCSE Staff. Deadlines and upcoming dates! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(1):10, January 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Staff:2019:DUDb

[SIG19b] SIGCSE Staff. Deadlines and upcoming dates! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(2):18, April 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Simon:2013:ACS

[Sim13] Simon. The Australasian chapter of SIGCSE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(4):6, October 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2017:KCI

[SM17] Judith Sheard and Calkin Suero Montero. Koli Calling International Computing Education Conference review. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(2):3, April 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sobel:2014:NSL

[Sob14] Ann E. K. Sobel. Nurturing student leadership skills. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(3):10, July 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sahami:2013:CSC

[SR13] Mehran Sahami and Steve Roach. Computer Science curricula 2013: getting involved and getting ready. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(2): 5–6, April 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2013:SSP

[SS13] Judy Sheard and Amber Settle. SIGCSE special project grants. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(4):5, October 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stephenson:2014:CRS

[Ste14] Chris Stephenson. CSTA resources for SIGCSE members. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(2):7, April 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sweedyk:2012:MS

[Swe12] Z. Sweedyk. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(4):7, October 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sweedyk:2013:ICU

[Swe13] Z. Sweedyk. ITiCSE 2013: Canterbury, UK. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(3):8, July 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tenenberg:2017:IPb

[TCM^{+}17] Josh Tenenberg, Donald Chinn, Lauri Malmi, Ari Korhonen, and Judy Sheard. ICER 2017 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(3):2, July 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tenenberg:2017:IPa

[TCSM17] Josh Tenenberg, Donald Chinn, Judy Sheard, and Lauri Malmi. ICER 2017 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 49(1):6, January 2017. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Townsend:2016:BPA

[Tow16] Gloria Childress Townsend. Broadening participation: add RE-SPECT to your summer plans. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 48(2):7, April 2016. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Turner:2013:CSL

[TY13a] Will Turner and Frank Young. Changes to the SIGCSE lists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(1):9, January 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Turner:2013:LSI

[TY13b] Will Turner and Frank Young. Life as SIGCSE information codirector: the error messages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(1):9, January 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tymann:2014:SCSa

[Tym14a] Paul Tymann. The 2014 SIGCSE Computer Science Education Awards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(1):4, January 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tymann:2014:NCE

[Tym14b] Paul Tymann. NSF computing education funding. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(1):8, January 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tymann:2014:SCSb

[Tym14c] Paul Tymann. SIGCSE computer science educator: awards nomination guide. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 46(3):8–9, July 2014. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tymann:2015:SCS

[Tym15] Paul Tymann. SIGCSE computer science education awards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 47(3):11, July 2015. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2010:APC

[Wal10] Henry Walker. The art of paper categorization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 42(3):5, September 2010. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2011:ACP

[Wal11a] Henry M. Walker. ACM copyrights and plagiarism. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 43(3):7–8, September 2011. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2011:PSS

[Wal11b] Henry M. Walker. Paper submissions for SIGCSE conferences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 43(4):6, October 2011. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2011:SNa

[Wal11c] Henry M. Walker. SIGCSE by the numbers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 43(1):4–5, March 2011. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2011:SNb

[Wal11d] Henry M. Walker. SIGCSE by the numbers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 43(2):7, June 2011. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2012:INC

[Wal12a] Henry M. Walker. ITiCSE 2012: networking and community in Haifa, Israel. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(3):3, July 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2012:SN

[Wal12b] Henry M. Walker. SIGCSE by the numbers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(2):6, April 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2012:SNJ

[Wal12c] Henry M. Walker. SIGCSE by the numbers: JavaScript. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(1):8, January 2012. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2013:MCC

[Wal13] Henry M. Walker. Mathematics with computing and computational science: dialog encouraged! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 45(2):7, April 2013. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wolz:2011:KCS

[Wol11] Ursula Wolz. K-12 Computer Science education news. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 43(4):3, October 2011. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zhang:2019:RFA

[ZCLR19] Ming Zhang, Stephen Cooper, and Andrew Luxton-Reilly. Report on the First ACM Global Computing Education Conference (CompEd). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(3):4–6, July 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zhang:2019:STS

[ZSH^{+}19] Jian Zhang, Mark Sherriff, Sarah Heckman, Pam Cutter, and Alvaro Monge. SIGCSE technical symposium 2020 call for submissions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 51(3):2–3, July 2019. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).